# Volumes of Regular Hyperbolic Simplices

### v1

@@ 1.1 Introduction
The hyperbolic volume of an ideal regular hyperbolic simplex $\Delta_{\infty}^{d}$ is known to be finite, for $d\geq 2$. The following explicit values are known, see [13], [23, p. 207] or [17, Equations (13.55),(13.56))]:
@@
$\operatorname{Vol}_{3,-1}(\Delta_{\ell}^{3})=\int_{0}^{\ell}\frac{3a\sinh a\,{\rm d}a}{(1+2\sinh a)\sqrt{(1+\cosh a)(1+3\cosh a)}}.$
@@ 4.5 Analytic continuation of an integral
By Corollary 4.5, $\Phi(\mu_{j}\sqrt{z}x)$ is bounded on $\{x\in\mathbb{C}:|\arg x|\leq\frac{\pi}{4}\}$. Hence, it suffices to show that $\int_{0}^{\pi/4}|{\rm e}^{-(A{\rm e}^{{\rm{i}}\,\mathrm{sgn}(\omega)\theta})^{2}/2}|{\rm d}\theta\to 0$ as $A\to+\infty$. For all $\theta$ in the interval of integration, we have
$\Big{|}{\rm e}^{-(A{\rm e}^{{\rm{i}}\,\mathrm{sgn}(\omega)\theta})^{2}/2}\Big{|}={\rm e}^{\operatorname{Re}(-(A{\rm e}^{{\rm{i}}\,\mathrm{sgn}(\omega)\theta})^{2}/2)}={\rm e}^{-\frac{A^{2}}{2}\operatorname{Re}(({\rm e}^{{\rm{i}}\,\mathrm{sgn}(\omega)\theta})^{2})}={\rm e}^{-\frac{A^{2}}{2}\cos(2\theta)}={\rm e}^{-\frac{A^{2}}{2}\sin(\frac{\pi}{2}-2\theta)}.$
@@
$\int_{0}^{\pi/4}|{\rm e}^{-(A{\rm e}^{{\rm{i}}\,\mathrm{sgn}(\omega)\theta})^{2}/2}|{\rm d}\theta=\int_{0}^{\pi/4}{\rm e}^{-\frac{A^{2}}{2}\sin(2x)}{\rm d}x.$
@@
$\int_{0}^{\pi/4}|{\rm e}^{-(A{\rm e}^{{\rm{i}}\,\mathrm{sgn}(\omega)\theta})^{2}/2}|{\rm d}\theta\leq\int_{0}^{\pi/4}{\rm e}^{-\frac{A^{2}}{2}cx}{\rm d}x=\frac{2}{cA^{2}}\left(1-{\rm e}^{-\frac{A^{2}}{2}c\frac{\pi}{4}}\right)\xrightarrow[A\to+\infty]{}0.$

### v2

@@ 1.1 Introduction
The hyperbolic volume of an ideal regular hyperbolic simplex $\Delta_{\infty}^{d}$ is known to be finite, for $d\geq 2$. The following explicit values are known, see [13], [23, p. 207] or [17, Equations (13.55),(13.56)]:
@@
$\operatorname{Vol}_{3,-1}(\Delta_{\ell}^{3})=\int_{0}^{\ell}\frac{3a\sinh a\,{\rm d}a}{(1+2\cosh a)\sqrt{(1+\cosh a)(1+3\cosh a)}}.$
@@ 4.5 Analytic continuation of an integral
By Corollary 4.5, $\Phi(\mu_{j}\sqrt{z}x)$ is bounded on $\{x\in\mathbb{C}:|\arg x|\leq\frac{\pi}{4}\}$. Hence, it suffices to show that $\int_{0}^{\pi/4}|{\rm e}^{-(A{\rm e}^{{\rm{i}}\,\delta\theta})^{2}/2}A\,{\rm{i}}\,\delta{\rm e}^{{\rm{i}}\,\delta\theta}|{\rm d}\theta\to 0$ as $A\to+\infty$. For all $\theta$ in the interval of integration, we have
$\Big|{\rm e}^{-(A{\rm e}^{{\rm{i}}\,\delta\theta})^{2}/2}A\,{\rm{i}}\,\delta{\rm e}^{{\rm{i}}\,\delta\theta}\Big|=A{\rm e}^{\operatorname{Re}(-(A{\rm e}^{{\rm{i}}\,\delta\theta})^{2}/2)}=A{\rm e}^{-\frac{A^{2}}{2}\operatorname{Re}(({\rm e}^{{\rm{i}}\,\delta\theta})^{2})}=A{\rm e}^{-\frac{A^{2}}{2}\cos(2\theta)}=A{\rm e}^{-\frac{A^{2}}{2}\sin(\frac{\pi}{2}-2\theta)}.$
@@
$\int_{0}^{\pi/4}|{\rm e}^{-(A{\rm e}^{{\rm{i}}\,\delta\theta})^{2}/2}A\,{\rm{i}}\,\delta{\rm e}^{{\rm{i}}\,\delta\theta}|{\rm d}\theta=\int_{0}^{\pi/4}A{\rm e}^{-\frac{A^{2}}{2}\sin(2x)}{\rm d}x.$
@@
$\int_{0}^{\pi/4}|{\rm e}^{-(A{\rm e}^{{\rm{i}}\,\delta\theta})^{2}/2}A\,{\rm{i}}\,\delta{\rm e}^{{\rm{i}}\,\delta\theta}|{\rm d}\theta\leq\int_{0}^{\pi/4}A{\rm e}^{-\frac{A^{2}}{2}cx}{\rm d}x=A\frac{2}{cA^{2}}\left(1-{\rm e}^{-\frac{A^{2}}{2}c\frac{\pi}{4}}\right)\xrightarrow[A\to+\infty]{}0.$